A helicopter flying with the wind made a 400 km trip in 2 hours and made a return trip against the wind in 2.5 hours. What was its speed in still air? What was the speed of the wind?
Answer by mananth(16949)

Helicopter speed =x mph
wind speed =y mph
against wind 2.5 hours
with wind 2 hours

Distance against 400 miles distance with 400 miles
t=d/r against wind -
400 / ( x - y )= 2.50
3 ( x - y ) = 400
2.50 x - 2.50 y = 400.00 ....................1

400.00 / ( x + y )= 2.00
2.00 ( x + y ) = 400.00
2.00 x + 2.00 y = 400.00 ...............2
Multiply (1) by 2.00
Multiply (2) by 2.50
we get y
5.00 x + -5.00 y = 800.00
5.00 x + 5.00 = 1,000.00
10.00 x = 1,800.00
/ 10.00
x = 180.00 mph

plug value of x in (1) y
2.50 x -2.50 y = 400.00
450.00 -2.50 -450.00 = 400.00
-2.50 y = 400.00
-2.50 y = -50.00 mph
y = 20.00
Helicopter speed 180.00 mph
wind speed 20.00 mph

m.ananth@hotmail.ca